How many liters of water vapor are produced when 23.6 g of NH4NO3 decomposes?
Express your answer with the appropriate units.

V=???
1
Expert's answer
2015-11-03T13:54:08-0500
2NH4NO3 = 2N2 + O2 + 4H2O
M(NH4NO3 ) = 14*2 + 1*4 + 16*3 = 80 (g/mol)
n (NH4NO3 ) = m/M = 23.6/80 = 0.295 (mol)
n(H2O) = 2* n(NH4NO3 ) = 0.59 (mol)
V(H2O) = n * Vm = 0.59 * 22.4 = 13.216 (l)

Answer
V (H2O) = 13.216 l